OpenStudy (anonymous):

Find the solution set for y = 2x + 1, given the replacement set {(-2, -3), (-1, -1), (0, 2), (1, 3), (2, 6)}.

jimthompson5910 (jim_thompson5910):

they basically want to know which points are on the line

jimthompson5910 (jim_thompson5910):

Let's check if (-2,-3) is on the line y = 2x + 1 -3 = 2*(-2) + 1 ... plug in x = -2 and y = -3 -3 = -4 + 1 -3 = -3 since the equation is true, this means (-2,-3) is on the line so part of the solution set is (-2,-3)

jimthompson5910 (jim_thompson5910):

do the same for the other points